OpenStudy (anonymous):

Could someone please help me with the derivative of this? Y= sqrt(4+3*f(x)) Thanks.

OpenStudy (neer2890):

you have to find \[ \frac{ dy }{ dx }???\]

OpenStudy (neer2890):

if you have to find dy/dx then \[y=\sqrt{4+3f(x)}\] \[\frac{ dy }{ dx }=\frac{ 1 }{ 2 }(4+3f(x))^{\frac{ -1 }{ 2 }}*3f \prime(x)\] =\[=\frac{ 3f \prime(x) }{ 2\sqrt{(4+3f(x))} }\]